# Fast Jet Tagging with MLP-Mixers on FPGAs

### v1

@@ 5 Results
In this section, we first evaluate the performance of the full-precision MLP-Mixer models on the jet tagging dataset and compare them with the JEDI-net models [9] using different numbers of input particles. We then analyse the performance, resource consumption, latency, and throughput of the deployable firmware for both MLP-Mixer and MLP models on the same dataset. Additionally, we compare quantized models with their full-precision counterparts and with previous studies to highlight the advantages of using MLP-Mixer-based architectures for this task. For fairness, the $p_{T}\geq 2$ GeV selection was not applied to the input particles when comparing our results with [22] and [9].
@@ 5.1 Full Precision Model
The performance of the full-precision MLP-Mixer models (MLPM-fp) and JEDI-net [9] models is summarized in Table 2. The four MLP-Mixer models take 16, 32, 64, and 128 particles as input, while the two most accurate JEDI-net variants use 100 and 150 input particles. To represent the input configuration, we use ${N_{p}}^{\mathrm{\#fea}}=\mathrm{\#ptl}$, where #$\mathrm{\#fea}$ denotes the number of features per particle and #$\mathrm{\#ptl}$ represents the maximum number of particles used as input. Both MLP-Mixer and JEDI-net models utilize all 16 particle features, as described in Section 3.1. For MLP-Mixer models, the top-$N$ particles are selected and ordered by $p_{T}$ before being fed into the model. In contrast, the JEDI-net models also take as input the top-$N$ particles by $p_{T}$ but do not rely on their specific order, as they utilize a permutation-invariant architecture.
The performance metrics, including the area-under-curve (AUC) and the true positive rate (TPR) at false positive rate (FPR) of $1\%$ and $10\%$, are presented in Table 2 for each class, with the best performance in each category highlighted in bold. The MLP-Mixer models consistently achieve higher AUC scores than the JEDI-net models across all classes. However, in terms of TPR at FPR=$1\%$ and FPR=$10\%$, their performance varies slightly, with some classes showing marginal improvements while others experience minor reductions. Among all MLP-Mixer variants, the model with 64 input particles delivers the best overall performance across all categories.
@@
Model JEDI-net [9] JEDI-net MLPM-fp MLPM-fp MLPM-fp MLPM-fp
w/ $\Sigma\mathcal{O}$ [9]
${N_{p}}^{16}=100$ ${N_{p}}^{16}=150$ ${N_{p}}^{16}=16$ ${N_{p}}^{16}=32$ ${N_{p}}^{16}=64$ ${N_{p}}^{16}=128$
AUC (%)
gluon 95.29 95.28 94.10 95.11 95.51 95.50
light quarks 93.01 92.90 92.20 93.09 93.40 93.35
W boson 97.39 96.95 96.09 97.51 97.84 97.83
Z boson 96.79 96.49 95.38 97.11 97.50 97.47
top quark 96.83 96.77 96.19 96.93 97.14 97.06
TPR at FPR=$10\%$
gluon 87.8 87.9 82.2 85.4 86.6 86.5
light quarks 82.2 81.8 78.0 80.3 81.0 81.2
W boson 93.8 92.7 89.5 93.1 93.9 93.9
Z boson 91.0 90.3 86.5 91.2 92.2 92.1
top quark 93.0 93.1 90.8 92.5 93.1 92.8
TPR at FPR=$1\%$
gluon 48.5 48.2 41.0 45.1 46.9 46.9
light quarks 30.2 30.1 25.8 28.5 28.2 28.1
W boson 70.4 65.8 51.8 70.4 75.4 75.3
Z boson 76.9 72.9 66.5 77.7 80.9 80.7
top quark 63.3 63.2 60.9 64.6 65.1 63.9
@@
22footnotetext: The reported FLOPs may be overestimated, as certain matrix multiplications involving constant one-hot matrices are included in the FLOP calculations. However, these operations can be efficiently implemented as gather or lookup operations, significantly reducing computational complexity, as noted in  [22].
@@
[…]
@@ 5.2 Quantized Models
Figure V: Accuracy vs. LUT usage and latency for the quantized MLP and JEDI-net models, each trained with 16 features per particle. Dashed lines indicate the accuracy of the corresponding full-precision models. Models marked with “$\times$” markers successfully underwent HDL synthesis but failed timing closure during the place & route phase.
Figures 4(a) and 4(b) illustrate the trade-off between accuracy, LUT usage, and latency for the quantized MLP-Mixer models, while Figure 5(a) and 5(b) present the same comparison for the quantized MLP models. Since our DA optimization involves explicit unrolling, both MLP-Mixer and MLP models operate without using any DSP slices. The dashed lines in the figures indicate the accuracy of the corresponding full-precision models with the same number of input particles. For comparison, we also show the LUT usage, latency, and accuracy of the quantized JEDI-net models for $N_{p}^{1}6$ = 30 and 50 from Ref. [22].
@@
Model ${N_{p}}^{16}$ Accuracy (%) Latency DSP LUT (k) FF (k) BRAM II
JEDI-net U4 [22] 50 80.9 130 (650 ns) 8,945 855 201 25 110
JEDI-net U5 [22] 50 81.2 181 (905 ns) 8,986 815 189 37 150
JEDI-net J4 [22] 32 78.4 58 (290 ns) 8,776 865 138 37 30
JEDI-net J5 [22] 32 79.9 181 (905 ns) 9,833 911 158 37 150
MLPMmax 16 77.5 15 (75 ns) 0 102 24 0 1
MLPMmax 32 80.7 14 (70 ns) 0 87 22 0 1
MLPMalt 32 79.9 13 (65 ns) 0 42 11 0 1
MLPMmax 64 81.6 14 (70 ns) 0 126 32 0 1
MLPMalt 64 81.3 13 (65 ns) 0 58 16 0 1
MLPMmax 128 81.3 16 (80 ns) 0 151 42 0 1
MLPmax 16 77.3 6 (30 ns) 0 73 18 0 1
MLPmax 32 79.7 7 (35 ns) 0 103 23 0 1
MLPmax 64 79.8 6 (30 ns) 0 57 15 0 1
MLPmax 128 80.0 7 (35 ns) 0 113 29 0 1
While the models saturate at different accuracy levels, we observe that the trade-off curves for the MLP-Mixer and MLP models with varying number of input particles remain mostly consistent – particularly in the accuracy region below 78% – with the exception of the 16-particles variants. This suggests that relying on parameter count or FLOPs as a proxy for resource consumption in quantized models can be misleading, as these metrics scale strongly with input size, as shown in Table 3. To provide further insights, Table 5 presents the bitwidths for weights and activations in the MLP-Mixer models, as defined in [11]. The data clearly shows that as the number of input particles increases, both the bitwidths and (1(1-( 1 -sparsity)))) of the weights and activations decrease, effectively compensating for the otherwise higher resource consumption.
@@
Model ${N_{p}}^{16}$ Weights Activations
Avg. BW Avg. BW Sparsity Avg. BW Avg. BW Sparsity
(non-zero) (non-zero)
MLPMmax 16 1.93 2.52 0.23 4.80 6.34 0.24
MLPMmax 32 1.55 2.51 0.38 3.27 5.66 0.42
MLPMalt 32 0.60 2.27 0.73 1.78 4.82 0.63
MLPMmax 64 1.06 2.44 0.57 2.55 5.19 0.51
MLPMalt 64 0.58 2.29 0.75 1.67 4.81 0.65
MLPMmax 128 0.57 2.27 0.75 1.78 4.76 0.62
@@ 5.4 Quantized model with $p_{T}$, $\eta$ and $\phi$ inputs
Figure VIII: Accuracy vs. LUT usage (left) and latency (right) of the quantized MLP-Mixer, Interaction Network (IN) [23], and Deep Sets (DS) [23] models trained using only $p_{T}$, $\eta$, and $\phi$ as input features per particle. Dashed lines indicate the accuracy of the corresponding full-precision models with the same number of input particles. Models marked with a “$\times$” successfully underwent HDL synthesis but failed timing closure during the place & route phase. The error bars for the IN and DS models reflect ambiguity in the reported accuracy in the original work [23].
@@
Figure IX: Accuracy vs. LUT usage (left) and latency (right) of the quantized MLP, Interaction Network (IN) [23], and Deep Sets (DS) [23] models trained with using only $p_{T}$, $\eta$, and $\phi$ as input features per particle. Dashed lines indicate the accuracy of the corresponding full-precision models with the same number of input particles. Models marked with a “$\times$” successfully underwent HDL synthesis but failed timing closure during the place & route phase. The error bars for the IN and DS models reflect ambiguity in the reported accuracy in the original work [23].
Table 6: Performance, latency, and resource consumption of the quantized MLP-Mixer, MLP, DS and IN models described by the accuracy and resource consumption. All models are trained on the hls4ml jet tagging dataset using only $p_{T}$, $\eta$, and $\phi$ as input features per particle. Only particles with $p_{T}\geq 2$ GeV are included in the training and evaluation. The intervals for accuracies for the DS and IN models are due to the ambiguity in the original work [23], where the author did not report model accuracies after quantization and firmware generation.
@@
Figure X: Receiver Operating Characteristic (RoC) curves of the MLPMmax models (solid lines) compared to their full precision counterpart (dashed lines) trained with 16, 32, 64, and 128 input particles and using only $p_{T}$, $\eta$, and $\phi$ as input features per particle. The Area Under the Curve (AUC) for each jet category is provided in the legend, formatted as “full-precision / quantized”.
In this section, we evaluate the performance of the quantized MLP-Mixer models trained using only $p_{T}$, $\eta$, and $\phi$ for each particle, rather than the full 16-feature set described in Table 1. This configuration reflects a more realistic experimental scenario as the derived kinematic features beyond momentum may not always be available in the hardware trigger systems. In addition to restricting input features, we apply a $p_{T}\geq 2$ GeV selection to each particle, as suggested in Ref. [23], to better approximate detector behavior. Apart from adjusting the layer sizes to accommodate the reduced feature set (as shown in Figure II with $n=3$), all other hyperparameters remain unchanged. Training and evaluation procedures follow the same methodology used for models trained with 16 features per particle, as detailed in Sections 5.1 and 5.2.
The performance of the quantized MLP-Mixer models trained with particle features $p_{T}$, $\eta$, and $\phi$ is shown in Figures 8(a) and 8(b). The corresponding results for the quantized MLP models we trained are shown in Figures 9(a) and 9(b). These models are compared with the 8-bits quantized Interaction Network (IN) and Deep Sets (DS) models proposed in Ref. [23] where they were trained with the same inputs. For the IN and DS models, the error bars on the accuracy reflect ambiguities in the reported performance in the original work, where only full-precision accuracy was provided. The accuracy of their quantized counterparts was estimated to be at least 90% of their full-precision versions. As shown in the figures, the quantized MLP-Mixer models consistently outperform the IN and DS models in both accuracy and resource consumption.
@@
Similar to models trained with 16 particles as input, the MLP models optimized with DA and HGQ achieve lower latency but are limited in accuracy compared to MLP-Mixer models. Nevertheless, our results demonstrate that when properly optimized, even MLP models can surpass permutation-invariant architectures (IN and DS models) in both resource efficiency and accuracy, as shown in Figure 9(a) and 9(b).
Finally, Figure X presents the ROC curves for the full-precision and best-performing quantized MLP-Mixer models trained with particle features $p_{T}$, $\eta$, and $\phi$. The quantized models achieve comparable, albeit slightly lower, performance relative to their full-precision counterparts.
@@ 6 Conclusions
Author information
Correspondence and material requests should be emailed to C. Sun (chsun@cern.ch).
@@
The data and software required to reproduce this work can be found at https://doi.org/10.5281/zenodo.3602260 [28] and https://github.com/calad0i/HGQ-demos/tree/master/jet_classifier_large.

### v2

@@ 5 Results
In this section, we first evaluate the performance of the full-precision MLP-Mixer models on the jet tagging dataset and compare them with the JEDI-net models [jedi-net] using different numbers of input particles. We then analyse the performance, resource consumption, latency, and throughput of the deployable firmware for both MLP-Mixer and MLP models on the same dataset. Additionally, we compare quantized models with their full-precision counterparts and with previous studies to highlight the advantages of using MLP-Mixer-based architectures for this task. For fairness, the $p_{T}\geq 2$ GeV selection was not applied to the input particles when comparing our results with [jedi-fpga] and [jedi-net].
@@ 5.1 Full Precision Model
The performance of the full-precision MLP-Mixer models (MLPM-fp) and JEDI-net [jedi-net] models is summarized in Table 2. The four MLP-Mixer models take 16, 32, 64, and 128 particles as input, while the two most accurate JEDI-net variants use 100 and 150 input particles. To represent the input configuration, we use ${N_{p}}^{\mathrm{\#fea}}=\mathrm{\#ptl}$, where $\mathrm{\#fea}$ denotes the number of features per particle and $\mathrm{\#ptl}$ represents the maximum number of particles used as input. Both MLP-Mixer and JEDI-net models utilize all 16 particle features, as described in Section 3.1. For MLP-Mixer models, the top-$N$ particles are selected and ordered by $p_{T}$ before being fed into the model. In contrast, the JEDI-net models also take as input the top-$N$ particles by $p_{T}$ but do not rely on their specific order, as they utilize a permutation-invariant architecture.
[…]
@@
[…]
@@
22footnotetext: The reported FLOPs may be overestimated, as certain matrix multiplications involving constant one-hot matrices are included in the FLOP calculations. However, these operations can be efficiently implemented as gather or lookup operations, significantly reducing computational complexity, as noted in  [jedi-fpga].
@@
JEDI-net [jedi-net] ${N_{p}}^{16}=100$ 33,625 116M5.1
JEDI-net w/ $\Sigma\mathcal{O}$ [jedi-net] ${N_{p}}^{16}=150$ 8,767 458M5.1
@@ 5.2 Quantized Models
Figure V: Accuracy vs. LUT usage and latency for the quantized MLP and JEDI-net models, each trained with 16 features per particle. Dashed lines indicate the accuracy of the corresponding full-precision models. Models marked with $\times$” markers successfully underwent HDL synthesis but failed timing closure during the place & route phase. These models’ LUT usage are can still be used for reference, but their latencies reported are not accurate.
Figures 4(a) and 4(b) illustrate the trade-off between accuracy, LUT usage, and latency for the quantized MLP-Mixer models, while Figure 5(a) and 5(b) present the same comparison for the quantized MLP models. Since our DA optimization involves explicit unrolling, both MLP-Mixer and MLP models operate without using any DSP slices. The dashed lines in the figures indicate the accuracy of the corresponding full-precision models with the same number of input particles. For comparison, we also show the LUT usage, latency, and accuracy of the quantized JEDI-net models for ${N_{p}}^{16}$ = 30 and 50 from Ref. [jedi-fpga].
@@
Model ${N_{p}}^{16}$ Accuracy (%) Latency DSP LUT (k) FF (k) BRAM II JEDI-net U4 [jedi-fpga] 50 80.9 130 (650 ns) 8,945 855 201 25 110 JEDI-net U5 [jedi-fpga] 50 81.2 181 (905 ns) 8,986 815 189 37 150 JEDI-net J4 [jedi-fpga] 32 78.4 58 (290 ns) 8,776 865 138 37 30 JEDI-net J5 [jedi-fpga] 32 79.9 181 (905 ns) 9,833 911 158 37 150 MLPMmax 16 77.5 15 (75 ns) 0 102 24 0 1 MLPMmax 32 80.7 14 (70 ns) 0 87 22 0 1 MLPMalt 32 79.9 13 (65 ns) 0 42 11 0 1 MLPMmax 64 81.6 14 (70 ns) 0 126 32 0 1 MLPMalt 64 81.3 13 (65 ns) 0 58 16 0 1 MLPMmax 128 81.3 16 (80 ns) 0 151 42 0 1 MLPmax 16 77.3 6 (30 ns) 0 73 18 0 1 MLPmax 32 79.7 7 (35 ns) 0 103 23 0 1 MLPmax 64 79.8 6 (30 ns) 0 57 15 0 1 MLPmax 128 80.0 7 (35 ns) 0 113 29 0 1
While the models saturate at different accuracy levels, we observe that the trade-off curves for the MLP-Mixer and MLP models with varying number of input particles remain mostly consistent – particularly in the accuracy region below 78% – with the exception of the 16-particles variants. This suggests that relying on parameter count or FLOPs as a proxy for resource consumption in quantized models can be misleading, as these metrics scale strongly with input size, as shown in Table 3. To provide further insights, Table 5 presents the bitwidths for weights and activations in the MLP-Mixer models, as defined in [hgq]. The data clearly shows that as the number of input particles increases, both the bitwidths and $(1$-sparsity)) of the weights and activations decrease, effectively compensating for the otherwise higher resource consumption.
@@
Model ${N_{p}}^{16}$ Weights Activations Avg. BW Avg. BW Sparsity Avg. BW Avg. BW Sparsity (non-zero) (non-zero) MLPMmax 16 1.93 2.52 0.23 4.80 6.34 0.24 MLPMmax 32 1.55 2.51 0.38 3.27 5.66 0.42 MLPMalt 32 0.60 2.27 0.73 1.78 4.82 0.63 MLPMmax 64 1.06 2.44 0.57 2.55 5.19 0.51 MLPMalt 64 0.58 2.29 0.75 1.67 4.81 0.65 MLPMmax 128 0.57 2.27 0.75 1.78 4.76 0.62
@@ 5.3 Feature Importance
5.4 Particle-wise Feature prioritization
In this section, we perform an ablation study to evaluate the impact of particle-wise feature prioritization on the efficiency of the MLP-Mixer models: different bitwidths are only assigned to the different channels, but all particles will share the same bitwidth for each channel. This is in contrast to the heterogeneous quantization used in Section 5.2, where each particle can have a different bitwidths, allowing the model to prioritize features based on their importance for each particle, as shown in Figure VII.
All other hyperparameters and other configuration remain unchanged from Section 5.2. The results are presented in Figure VIII for the accuracy vs. LUT usage trade-off. Latency vs. accuracy trade-off is not shown as the majority of models do not meet timing, and the latency reported this way will not be accurate. As the models are fully and pipelined, adding more pipeline stages to close timing would not incur significant LUT overhead, and therefore the models LUT usage with failed timing are still useful for comparison.
Comparing to the results shown in Figure 4(a), the models with uniform bitwidths for each input particle require significant more LUTs to achieve the same accuracy as the models without it. The difference is particularly pronounced for the models with 64 and 128 input particles, where the models with uniform particle-wise quantization require one order of magnitude more LUTs to achieve the same accuracy as the models with fully heterogeneous quantization. As uniform bitwdiths for each particle is a prerequisite for permutation invariance neural networks, these results strongly suggests that permutation invariance could lead to major resource overhead.
Figure VIII: Accuracy vs. LUT usage and latency for the quantized MLP-Mixer without particle-wise heterogeneous quantization, each trained with 16 features per particle. Dashed lines indicate the accuracy of the corresponding full-precision models. Models marked with $\times$” markers successfully underwent HDL synthesis but failed timing closure during the place & route phase. These models’ LUT usage are can still be used for reference, but their latencies reported are not accurate.
@@ 5.5 Quantized model with $p_{T}$, $\eta$ and $\phi$ inputs
Figure IX: Accuracy vs. LUT usage (left) and latency (right) of the quantized MLP-Mixer, Interaction Network (IN) [ds-fpga], and Deep Sets (DS) [ds-fpga] models trained using only $p_{T}$, $\eta$, and $\phi$ as input features per particle. Dashed lines indicate the accuracy of the corresponding full-precision models with the same number of input particles. Models marked with a $\times$” successfully underwent HDL synthesis but failed timing closure during the place & route phase. These models’ LUT usage are can still be used for reference, but their latencies reported are not accurate. The error bars for the IN and DS models reflect ambiguity in the reported accuracy in the original work [ds-fpga].
@@
Figure X: Accuracy vs. LUT usage (left) and latency (right) of the quantized MLP, Interaction Network (IN) [ds-fpga], and Deep Sets (DS) [ds-fpga] models trained with using only $p_{T}$, $\eta$, and $\phi$ as input features per particle. Dashed lines indicate the accuracy of the corresponding full-precision models with the same number of input particles. Models marked with a $\times$” successfully underwent HDL synthesis but failed timing closure during the place & route phase. These models’ LUT usage are can still be used for reference, but their latencies reported are not accurate. The error bars for the IN and DS models reflect ambiguity in the reported accuracy in the original work [ds-fpga].
Table 6: Performance, latency, and resource consumption of the quantized MLP-Mixer, MLP, DS and IN models described by the accuracy and resource consumption. All models are trained on the hls4ml jet tagging dataset using only $p_{T}$, $\eta$, and $\phi$ as input features per particle. Only particles with $p_{T}\geq 2$ GeV are included in the training and evaluation. The intervals for accuracies for the DS and IN models are due to the ambiguity in the original work [ds-fpga], where the author did not report model accuracies after quantization and firmware generation.
@@
Figure XI: Receiver Operating Characteristic (RoC) curves of the MLPMmax models (solid lines) compared to their full precision counterpart (dashed lines) trained with 16, 32, 64, and 128 input particles and using only $p_{T}$, $\eta$, and $\phi$ as input features per particle. The Area Under the Curve (AUC) for each jet category is provided in the legend, formatted as “full-precision / quantized”.
In this section, we evaluate the performance of the quantized MLP-Mixer models trained using only $p_{T}$, $\eta$, and $\phi$ for each particle, rather than the full 16-feature set described in Table 1. This configuration reflects the minimal set of features may be provided by the hardware trigger systems, where all kinematic features beyond momentum may not always be available at L1T. In addition to restricting input features, we apply a $p_{T}\geq 2$ GeV selection to each particle as performed in Ref. [ds-fpga], to better approximate detector behavior. Apart from adjusting the layer sizes to accommodate the reduced feature set (as shown in Figure II with $n=3$), all other hyperparameters remain unchanged. Training and evaluation procedures follow the same methodology used for models trained with 16 features per particle, as detailed in Sections 5.1 and 5.2.
The performance of the quantized MLP-Mixer models trained with particle features $p_{T}$, $\eta$, and $\phi$ is shown in Figures 9(a) and 9(b). The corresponding results for the quantized MLP models we trained are shown in Figures 10(a) and 10(b). These models are compared with the 8-bits quantized Interaction Network (IN) and Deep Sets (DS) models proposed in Ref. [ds-fpga] where they were trained with the same inputs. For the IN and DS models, the error bars on the accuracy reflect ambiguities in the reported performance in the original work, where only full-precision accuracy was provided. The accuracy of their quantized counterparts was estimated to be at least 90% of their full-precision versions. As shown in the figures, the quantized MLP-Mixer models consistently outperform the IN and DS models in both accuracy and resource consumption.
@@
Similar to models trained with 16 particles as input, the MLP models optimized with DA and HGQ achieve lower latency but are limited in accuracy compared to MLP-Mixer models. Nevertheless, our results demonstrate that when properly optimized, even MLP models can surpass permutation-invariant architectures (IN and DS models) in both resource efficiency and accuracy, as shown in Figure 10(a) and 10(b).
Finally, Figure XI presents the ROC curves for the full-precision and best-performing quantized MLP-Mixer models trained with particle features $p_{T}$, $\eta$, and $\phi$. The quantized models achieve comparable, albeit slightly lower, performance relative to their full-precision counterparts.
@@ 6 Conclusions
For future works, we identify a few improvements could be made over this work, and hope they could be useful for the community:
  • the kernels for the particle-wise mixers (MLP2 and MLP4 in Figure II) could be regularized for mitigating overfitting. In the models discussed in this work, only the ${N_{p}}^{16}=128$ models showed slight signs of overfitting in the full-precision and a few least quantized models. However, this could be more pronounced with larger models or datasets.
  • In this work, for the feature-wise mixers (MLP1 and MLP3), the kernel applied on each particle is exactly identical. It may be beneficial to allow for heterogeneous bitwidths the kernels applied to each particle to further optimize resource usage. This would require modifications in the quantization frameworks, but the authors believe that the benefits would be worth the effort.
  • For distributed arithmetic implementation, this work uses HLS for implementing the firmware. Targeting specific FPGA architecture, direct mapping the logic to LUTs and other logic primitives may be beneficial. For instance, using a compressor trees for summing over the partial outputs of the distributed arithmetic would lead to further reduction in resource usage.
@@
The data and software required to reproduce this work can be found at https://doi.org/10.5281/zenodo.3602260 [hls4ml-dataset] and https://github.com/calad0i/HGQ-demos/tree/master/jet_classifier_large.